what is the mass of a sphere of diameter 1cm whose density is 900 kg/m^3

Grade:

7 Answers

Prajwal kr
11 years ago

radius=0.5cm

Volume=4/3pi x (0.5)3

          =0.52 cm3

          = 0.52 x 10-6 m3

 

Mass= Density x Volume

        = 900 x 0.52x10-6  kg

        =0.468 g

lokesh soni
11 years ago

radius=.5

volume=4∏r3/3

v=5.2333333333334*10-7m3

mass=density *volume

mass=

5.2333333333334*10-7m3*900kg

0.000471kg


VEERU YERRAYYA
11 years ago

density = mass/ volume

given denstiy = 900kg/m3 and diameter = 1cm so radius(R) = 0.5 cm = 0.005m

900 = mass/ (4/3 π R3 )

Mass = (900*3 π R3 )/4 = (900*3* π *(0.005)3 )/ 4 = 2.650718801×10?4 Kg

naninagaraju sivacharan lakshmi narsing
11 years ago

given diameter of the sphere=1cm

so that the radius is=diameter/2

=1/2=0.5cm

density of the sphere=900kg/m^3

formula: density=mass/volume

volume of the sphere =4/3 pi r^3

where r=radius of the sphere

mass=density*volume

      ==>900*4/3*3.14*(0.5*10^-2m)^3

     ===>4.71*10^-4

Thanks for asking these type of questions

shanmukha rella
11 years ago

mass of the sphere is = 4.713 × 10^−4 kg

Khandavalli Satya Srikanth
11 years ago

mass=volume * density.By using this formula we can find the mass of sphere.

Volume of sphere=(4/3)pi*r^3

V=(4/3)*3.14*(1*10^-2)^3=4.187*10^-6

Mass=(4.187*10^-6)*900=3768*10^-6kg

JAYANT ARORA
11 years ago

 Mass = 3.77 x 10-3 kg..